The rate at which a certain drug is eliminated by the body follows first-order kinetics, with a half life of 79 minutes.

Suppose in a particular patient the concentration of this drug in the bloodstream immediately after injection is

0.20/μgmL. What will the concentration be 395 minutes later?

Round your answer to 2 significant digits.
Answer #1

number of half life = total time / half life

= 395 / 79

= 5

After n half lifes:

N/No = (1/2)^n

N / 0.20 = (1/2)^5

N / 0.20 = 1/32

N = 0.20/32

N = 6.25*10^-3 /ug/mL

Answer: 6.3*10^-3 /ug/mL